There are 4 red, 8 yellow, and 6 blue socks in a drawer. Once a sock is selected, it isn't replaced. What is the probability of reaching in the drawer without looking & choosing 2 blue socks?
1
Expert's answer
2012-02-10T07:55:07-0500
There are 4+8+6 = 18 socks in a drawer at all. The probability of choosing blue socks is Pb = 6/18 = 1/3, cause there are 6 blue socks among 18 different socks. The probability of choosing two blue socks without replacing is Pbb = Pb² = (1/3)² = 1/9.